### v2

# COCOLA: Coherence-Oriented Contrastive Learning of Musical Audio Representations

Abstract
We present COCOLA (Coherence-Oriented Contrastive Learning for Audio), a contrastive learning method for musical audio representations that captures the harmonic and rhythmic coherence between samples. Our method operates at the level of stems (or their combinations) composing music tracks and allows the objective evaluation of compositional models for music in the task of accompaniment generation. We also introduce a new baseline for compositional music generation called CompoNet, based on ControlNet, generalizing the tasks of MSDM, and quantify it against the latter using COCOLA. We release all models trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
1 Introduction
Recently, there have been significant advances in music generation in the continuous domain [1, 2, 3, 4, 5], thanks to the impressive development of generative models [6, 7, 8]. In addition to producing high-quality tracks of increasing length [5], these models offer precise semantic control through textual conditioning [9, 10]. However, they are limited as tools for musical composition, since they output a final mix containing all stems. To overcome this, a new range of compositional generative models is emerging [11, 12, 13], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive. The most important application of these models is accompaniment generation, where, given multiple conditioning sources (combined or not), the model is asked to output a new set (or a mixture) of coherent stems. Although previous models could generate accompaniments [14, 15], they could not be used iteratively (acting sequentially on the model’s outputs) in a composition process.
A significant problem with this line of research is the lack of an objective metric for quantifying the coherence of the generated outputs w.r.t. the inputs. For example, [11] proposes the sub-FAD metric as a multi-stem generalization of the FAD [16] protocol proposed in [15]. However, this metric is not optimal for assessing coherence, as it focuses on global quality instead of the level of harmony and rhythm shared by constituent stems.
Figure 1: Illustration of COCOLA score. COCOLA is a contrastive model able to estimate the coherence between instrumental tracks and generated accompaniments.
To this end, we propose a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1). The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub-mixtures of stems) and minimizing it on sub-components belonging to different windows. With the model, we define a COCOLA score as the similarity between conditioning tracks and accompaniments in the embedding space.
Additionally, given the scarcity of open-source compositional music models (to our knowledge, only MSDM is available publicly [11]), we introduce and release [^1] https://github.com/gladia-research-group/cocola an improved latent diffusion model called CompoNet, based on ControlNet [17], which generalizes the tasks of previous compositional models. We benchmark CompoNet against MSDM with the COCOLA score and FAD [16] on accompaniment generation, showcasing better performance.
After discussing related work in Section 2, we introduce COCOLA and CompoNet in Section 3. We describe the experimental setup in Section 4 and present the results in Section 5. We conclude the article in Section 6.
Figure 2: The COCOLA training procedure (single stem case). We first randomly crop windows of size $L$ from a batch of $K$ tracks (depicted on the left). As a second step, we randomly select two distinct stems in each window. For example, in the first window we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums). Thus, we embed all selected stems with the COCOLA encoder $f_{\theta}$, obtaining latent representations. For example, we obtain $\mathbf{h}^{1}_{1}$ and $\mathbf{h}^{1}_{2}$ from the first window. Finally, we compute the contrastive loss (Eq. (3)) considering embedings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
2 Related Work
2.1 Contrastive Methods for Audio
Contrastive learning [18, 19] can be formulated both as a supervised or self-supervised problem.
Supervised contrastive learning methods are typically cross-modal, requiring labeled information alongside audio data. In early works, the labeled information was in the form of simple tags, while the loss used to align embeddings of audio segments and tags was the triplet loss [20]. Within the same data setting,[21] used the contrastive loss of SimCLR [22]. With the advent of the transformer architecture [23], using complex sentences instead of simple tags became feasible. MuLaP [24] is the first model to train a common representation between audio and sentences in the musical domain. In such work, the audio and text are processed by a joint transformer encoder, conveying information about the two modalities through cross-attention layers. Although it is not a contrastive model per se, an audio-text matching loss uses negative examples to encourage the model to focus on aligned pairs. More recent works [25, 26, 9, 27], consider separate textual and audio encoders, which makes it possible to use the two branches independently at inference time.
Self-supervised representation learning methods [28, 29, 30, 31] build embedding spaces targeting structural information extracted from the audio data itself. In [32], the authors build positive examples for a triplet loss by augmenting with Gaussian noise, time and frequency translations, and sampling with time proximity. They also consider example mixing. While we compare coherent mixes in our method (Section 3.1), in [32], positive pairs are not coherence-related (e.g., mixing siren and dog sounds). As in the supervised case, following [22], multi-class cross-entropy losses are employed [33, 34, 35]. In COLA [33], the authors train an embedding model with contrastive loss using the simple criterion of sampling positive pairs only from the same audio track (still employing Gaussian noise), outperforming a fully supervised baseline in a plethora of tasks. [36] pairs mixtures with sources extracted via source separation.
The proposed COCOLA method shares aspects of both supervised and self-supervised approaches. Given that stems are pre-separated, we cannot consider the method purely self-supervised. At the same time, we process such data with a uni-modal encoder, as is the case of self-supervised methods.
2.2 Compositional Waveform Music Generation
Compositional music generation in the waveform domain (as opposed to symbolic domains such as MIDI or sheet music [37]) was introduced by [11], proposing the Multi-Source Diffusion Model (MSDM). Such a model captures the joint distribution of a fixed set of sources (e.g., Bass, Drums, Guitar, and Piano) using a (score-based) diffusion model [38, 39, 7, 6, 40]. At inference time, it is possible to perform unconditional generation of all sources, accompaniment generation of a subset given the complement, and source separation.
Following this initial work, AUDIT [41] proposes a diffusion model conditioned by a T5 Encoder [10], trained with instructions that allow the addition, removal (drop), and replacement of sources in an input audio mixture. This model operates on general audio signals with weak dependencies between the sources (e.g., environmental sounds). While MSDM is an unconditional generative model that processes single sources in parallel, AUDIT is a conditional generative model that processes mixtures sequentially.
InstructME [12] introduces AUDIT in the musical setting of MSDM, where sources are highly interdependent. Besides conditioning on chords, the fundamental difference lies in how the audio input is provided to the model: while in AUDIT, the input and output are two channels of the tensor that the diffusion model processes (the input is inpainted at inference time, similarly to how accompaniment generation is performed in MSDM), in InstructME, the input is processed by a convolutional network that follows the structure of the diffusion model U-Net [42] encoder. The features processed by the latter are aggregated with the features of the U-Net encoder (authors do not specify if they sum or inject the features).
StemGen [13] generates a single accompaniment source, given an instrument tag and an input audio mixture, via a masked music language model [4].
Another line of research focuses on inference-only methods for compositional music tasks, given pre-trained generative models. Based on generative source separation via Bayesian inference [43, 44, 45], Generalized Multi-Source Diffusion Inference (GMSDI) [46] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
Our proposed CompoNet baseline (Section 3.3) is a sequential conditional model that, unlike previous models, can perform all compositional music generation tasks (Figure 3). Differently than InstructME, the model is conditioned via a ControlNet adapter [17], which enables fine-tuning of diffusion models pre-trained with a large quantity of mixtures. Table 1 summarizes a comparison between CompoNet and current music compositional models.
3 Method
3.1 Stem-Level Contrastive Learning
In our setting, we have access to a dataset $D=\{\bar{\mathbf{x}}^{k}\}_{k=1,\dots,\bar{K}}$ containing $\bar{K}$ musical tracks $\bar{\mathbf{x}}^{k}$, each separated into a variable number $N$ of individual stems $\bar{\mathbf{x}}^{k}_{n}$, i.e., $\bar{\mathbf{x}}^{k}=\{\bar{\mathbf{x}}^{k}_{n}\}_{n=1,\dots,N}$. As a first step, we sample a batch of $K<\bar{K}$ tracks $\{\bar{\mathbf{x}}^{k}\}_{k=1,\dots,K}$ from $D$, with possible repetitions. Following, we slice a window $\mathbf{x}^{k}$ of size $L$ for each track $\bar{\mathbf{x}}^{k}$ in the batch (all stems in a window share the same length), such that no window contained in the same track overlaps for more than a ratio $r$, obtaining a new batch $\{\mathbf{x}^{k}\}_{k=1,\dots,K}$. Afterward, we select, for each $k$, two disjoint non-empty stem subsets $X_{1}^{k},X_{2}^{k}$ of $\mathbf{x}^{k}$. We define the sub-mixes $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ by summing the stems in $X_{1}^{k},X_{2}^{k}$:
$\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}$ (1)
When $X_{1}^{k},X_{2}^{k}$ are singletons, the sub-mixes are simply two stems in the window (single stem case). We work with sub-mixes because current compositional music generation methods [13] operate over them, including our proposed CompoNet (see Section 3.3). Like in COLA[33], we use a convolutional audio-only encoder [^2] In our notation, we incorporate into $f_{\theta}$ any domain transform preceding or following the convolutional network operations, like the (pre) mel-filterbank map and the (post) projection head $g$ in COLA. $f_{\theta}:\mathbb{R}^{L}\to\mathbb{R}^{d}$, mapping $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ to lower-dimensional embedding vectors $\mathbf{h}^{k}_{1}=f_{\theta}(\mathbf{m}^{k}_{1})$ and $\mathbf{h}^{k}_{2}=f_{\theta}(\mathbf{m}^{k}_{2})$, with $d$ the embedding dimension.
The COCOLA training procedure maximizes the agreement between pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}$ of sub-mixes embeddings in the same window. It decreases it for pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}$ ($j\neq k$) of sub-mixes embeddings in different windows. As in COLA, we use a bilinear similarity metric:
$\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})=(\mathbf{h}^{k}_{1})^{T}\mathbf{W}\mathbf{h}^{j}_{2}\,,$ (2)
where $\mathbf{W}$ is a learnable matrix. The loss we optimize is the multi-class cross entropy:
$\mathcal{L}=-\sum^{K}_{k=1}\log\frac{\exp(\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}))}{\sum^{K}_{j=1}\exp(\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}))}\,.$ (3)
We depict the training procedure of COCOLA in Figure 2 for the single stem case.
Figure 3: Inter-stem compositional generation tasks. $Y$ and $X$ represent the input and output stem sub-sets, respectively.
In the COLA training procedure, the positive pairs are (fully mixed) windows belonging to the same track. In COCOLA, they are sub-mixes belonging to the same window. As such, we allow for negative pairs belonging to the same track but in different windows. The $r$ ratio has to be chosen well to avoid strong overlaps between windows in the same track. In that case, we could potentially consider (nearly) coherent sub-mixes as negative pairs.
Model Task Methodology Input Output Coherence
UG AG SS EA ER
MSDM [11] ✓(MS) ✓(MS) Training / Supervised Multi / Source Multi / Source
GMSDI [46] ✓(MS) ✓(MS) Inference / Weakly Supervised Multi / Sub-mix Multi / Sub-mix
StemGen [13] ✓(1S) Training / Supervised Single / Sub-mix Single / Source
Audit [41] ✓(Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
InstructME [12] ✓(Extract 1S; Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
[…]
Table 1: Compositional audio models comparison. The various tasks are illustrated in Figure 3. 1S vs MS: the task operates on one vs multiple sources at a time. Multi vs Single on Input / Output: the model accepts multiple vs single inputs / outputs. Source vs Sub-mix on Input / Output: the model processes single sources or sub-mixes as inputs / outputs.
3.2 COCOLA Score
Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output). The model’s variables can be either a set of stems or sub-mixes. Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$. We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems). We define the COCOLA score (CCS) between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as:
$\text{CCS}(\mathbf{y},\tilde{\mathbf{x}})=\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})\,,$ (4)
the similarity (Eq. (2)) between their embeddings. The described procedure is depicted in Figure 1.
3.3 CompoNet
In order to showcase the utility of the COCOLA score as a metric for measuring the coherence of generated accompaniments, we propose an improved compositional model for music called CompoNet and compare it with the MSDM baseline in Section 5.
CompoNet first pre-trains a latent diffusion model $\epsilon_{\phi}$, based on U-Net architecture [42] conditioned via cross-attention layers [23], on a large dataset of tuples $(\mathbf{m},\mathbf{c})$ comprising audio mixtures $\mathbf{m}$ and relative textual descriptions $\mathbf{c}$. The mixtures $\mathbf{m}$ are mapped to latent vectors $\mathbf{z}=E^{\text{VAE}}(\mathbf{m})$ using a pre-trained VAE encoder [47], while the text descriptions are mapped to a continuous sequence $\mathbf{s}=E^{\text{TXT}}(\mathbf{c})$ using a text encoder (e.g., [10]). Following the DDPM formulation [7], the model is trained to reverse the forward Gaussian noising process given by:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\,,$ (5)
where $t\in[0,T]$ is a time index (with $T$ the maximum time step), $\bar{\alpha}_{t}$ is defined by integrating a noise schedule [7], and $\boldsymbol{\epsilon}$ is sampled from the standard Gaussian distribution. The model is trained with denoising score matching [39]:
$\min_{\phi}\mathbb{E}_{\mathbf{z},\boldsymbol{\epsilon},\mathbf{s},t}\left[\|\boldsymbol{\epsilon}-\epsilon_{\phi}(\mathbf{z}_{t},\mathbf{s},t)\|^{2}_{2}\right]\,,$ (6)
with $\mathbf{z}_{t}$ obtained via Eq. (5).
In a second phase, we fine-tune a ControlNet [17] adapter $C_{\psi}$ for tackling all compositional musical tasks (Figure 3) with a single model. The U-Net $\epsilon_{\phi}$ comprises an encoder, bottleneck, decoder structure $\epsilon_{\phi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E}}$. The ControlNet adapter is defined as $C_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E_{\phi_{E}}(\mathbf{z}_{t}+\text{conv}_{\text{in}}(\mathbf{w}),\mathbf{s},t)$, where $\text{conv}_{\text{in}}$ is a zero-initialized convolutional layer, and $\mathbf{w}$ is a latent (VAE) embedding of an external audio input. The ControlNet adapter outputs the set of processed features $\{C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})\}_{i=1,\dots,I}$ for each layer of $E_{\phi_{E}}$, with $I$ the total number of layers. The full ControlNet conditional architecture is defined as $\epsilon_{\phi,\psi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E},\psi}$ with $E_{\phi_{E},\psi}$ combining the encoder and ControlNet adapter features at each layer $i$ with zero-initialized convolutional layers $\text{conv}_{i}$:
$E^{i}_{\phi_{E},\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E^{i}_{\phi_{E}}(\mathbf{z}_{t},\mathbf{s},t)+\text{conv}_{i}(C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w}))\,.$
While we have described the general ControlNet architecture, we still have to describe how we train it in CompoNet, namely, the roles of the $\mathbf{z},\mathbf{w}$ and $\mathbf{s}$ variables. Iterating over a dataset containing tuples $(\mathbf{x},\mathbf{t})$ with multi-stem tracks $\mathbf{x}=\{\mathbf{x}_{n}\}_{n=1,\dots,N}$ and tag descriptions $\mathbf{t}=\{\mathbf{t}_{n}\}_{n=1,\dots,N}$ for each stem, we sample from $\mathbf{x}$ two arbitrary subsets of stems $Y,X\subseteq\mathbf{x}$, with $|X|>0$. $Y$ contains input stems while $X$ contains output stems. The topological relationships between such subsets define all possible compositional tasks, as depicted in Figure 3. While previous models partially solve some tasks (see Table 1), ours is the first to solve all of them simultaneously. We proceed like in Eq. (1) and mix the sources in $Y$ and $X$, obtaining $\mathbf{m}_{Y}$ and $\mathbf{m}_{X}$, respectively. Afterward, we encode them in the VAE latent space, defining $\mathbf{z}=E^{\text{VAE}}(\mathbf{m}_{X})$ and $\mathbf{w}=E^{\text{VAE}}(\mathbf{m}_{Y})$. We define the following prompt $\mathbf{s}$:
$\mathbf{s}=E^{\text{TXT}}(\mathbf{t}_{Y_{1}},\dots,\mathbf{t}_{Y_{|Y|}},\verb|SEP|,\ \mathbf{t}_{X_{1}},\dots,\mathbf{t}_{X_{|X|}})\,,$ (7)
specifying input and output mixture tags separated by a special token SEP.
Having specified the required inputs and outputs, we train $\epsilon_{\phi,\psi}$ via Eq. (6), optimizing only the $\psi$ weights. The $\mathbf{s}$ prompt instructs the model which task to perform based on the specified stem tags. The user knows such information during the composition process in a generative DAW [48]. While we only test CompoNet on accompaniment generation against MSDM using COCOLA, we want the models to share similar abilities for a fair comparison.
4 Experimental Setup
Test Dataset
Train Dataset MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
MoisesDB [49] 52.56 53.01 51.22 60.32
Slakh2100 [50] 53.06 53.58 53.78 59.35
CocoChorales [51] 70.10 61.48 67.50 99.78
All 90.43 93.06 90.06 99.89
Table 2: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
4.1 Datasets
In our experiments, we use four different stem-separated public datasets for training COCOLA and fine-tuning CompoNet. The datasets are MUSDB18-HQ [52], MoisesDB [49], Slakh2100 [50] and CocoChorales [51].
MUSDB18-HQ [52] is the uncompressed version (in WAV format) of the MUSDB18 dataset, initially introduced in [53]. This dataset is a standard for evaluating music source separation systems. It comprises 150 tracks – 100 for training and 50 for testing – totaling approximately 10 hours of professional-quality audio. Each track is divided into four stems: Bass, Drums, Vocals, and Other, with “Other” covering any components not classified under the first three categories.
MoisesDB [49] features 240 music tracks across diverse genres and artists, accumulating more than 14 hours of music. Unlike MUSDB18-HQ, MoisesDB is a genuine multi-track dataset, offering a two-tier taxonomy of 11 distinct stems. Each stem in this dataset includes more detailed type annotations (e.g., Guitar might be labeled as Acoustic Guitar or Electric Guitar). Not having pre-computed splits, we set a custom 0.8 (train) / 0.1 (validation) / 0.1 (test) split.
Slakh2100 [50] is synthesized from the Lakh MIDI Dataset v0.1 [54] employing high-quality sample-based virtual instruments. It features 2100 tracks organized into 1500 tracks for training, 375 for validation, and 225 for testing, together amounting to 145h of audio. The tracks are annotated into 34 stem categories. While such a dataset contains an order of magnitude more data than MUSDB18-HQ and MoisesDB, it does not share the same level of realism as the latter, being the tracks synthesized from MIDI.
CocoChorales [51] is a chorale audio music dataset created through a synthesis process like Slakh2100. However, it comprises a substantially vaster collection of 240000 tracks, extending over 1411 hours of audio data. It is produced by generating symbolic notes via a Coconet model, performing their synthesis with MIDI-DDSP [55]. This dataset is richly annotated, featuring details on performance attributes and synthesis parameters. CocoChorales includes a diverse range of 13 instruments spanning Strings, Brass, Woodwind, and Random ensembles. In our experiments we use the tiny version, comprising 4000 tracks.
4.2 Model Implementation
To implement the COCOLA encoder $f_{\theta}$, we follow [33] and employ the EfficientNet-B0 [56] convolutional architecture followed by a linear projection layer, operating on the mel-filterbank audio representation. The embedding dimension is 512. With respect to the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
For CompoNet, we employ the AudioLDM2 [57, 58] architecture. Since the authors pre-train the model on a large array of datasets [59, 60, 61, 62, 63, 64, 65, 66], we skip the pre-training phase in Section 3.2 and directly fine-tune a ControlNet adapter based on the AudioLDM 2-Large checkpoint [^3] https://huggingface.co/cvssp/audioldm2-large. During fine-tuning, we directly pass the conditioning prompt in Eq. 7 to the text-embedding mechanism of AudioLDM2 (based on CLAP [9, 27], T5 [10], and GPT2 [67]), conditioning both the U-Net and the ControlNet adapter.
100100100100200200200200300300300300400400400400500500500500600600600600700700700700800800800800000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Training stepTrain accuracyCosine similarityBilinear similarity
Figure 4: Cosine vs bilinear similarity during training. Training is performed with “COCOLA All”. The accuracy metric is defined in Eq. (8). We adopt the bilinear similarity in Eq. (2), given its improved performance.
Method
FAD $\downarrow$
CLAP
FAD $\downarrow$
EnCodec
FAD $\downarrow$
VGGish
COCOLA score $\uparrow$
All
COCOLA score $\uparrow$
CocoChorales
Slakh2100
MSDM [11] 0.23 92.81 2.01 3.31 0.72
CompoNet 0.30 106.23 3.20 13.50 8.32
Random 0.064 51.44 0.16 0.069 1.06
[…]
MUSDB18-HQ
MSDM [11] 0.29 148.09 2.36 11.61 3.37
CompoNet 0.37 130.04 2.14 11.94 7.21
Random 0.11 100.25 0.35 4.40 3.47
[…]
Table 3: Comparison between MSDM and CompoNet.
4.3 Training Details
All COCOLA models are trained on an NVIDIA RTX 4070 Super with 12GB of VRAM. Each training batch contains 32 audio chunks of 5s (16kHz). We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer [68] with a $10^{-3}$ learning rate. We add Gaussian noise to positive samples as a data augmentation method, with $\sigma=10^{-3}$. We tried training with cosine similarity [22] but we reported (Figure 4) lower performance, corroborating [33].
CompoNet models are also fine-tuned on NVIDIA GPUs (RTX 4080 16GB on MUSDB18-HQ; A10G, 24GB on Slakh2100). The training batches have size 2 (MUSDB18-HQ) and 5 (Slakh2100) and contain windows of 10.24s (16kHz). We train with Adam [68] with a $10^{-4}$ learning rate. We do not use the empty prompt for classifier-free guidance [69] during fine-tuning.
5 Experiments
We employ four COCOLA encoder models in our experiments : “COCOLA MoisesDB”, “COCOLA Slakh2100”, “COCOLA CocoChorales” and “COCOLA All”. The first three are trained on the homonym datasets, while the last one is trained on all three combined. For the “COCOLA CocoChorales” we use all ensables while on “COCOLA All” we use only the Random ensamble for a more balanced partitioning, with respect to the other datasets. MUSDB18-HQ, being the smallest dataset, is used as a held-out test dataset for studying generalization.
5.1 Coherent Sub-Mix Classification
We cross-test the performance of all COCOLA models, performing classification of coherent pairs on the test split of our datasets. More specifically, given an encoder $f_{\theta}$, we iterate a test set, collecting at each step a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$. Following the steps in Section 3.1 we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$. We define the accuracy over a batch as:
$\frac{1}{K}\sum_{k=1}^{K}\mathds{1}\biggl{(}k=\operatorname*{arg\,max}_{j\in[K]}\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})\biggr{)}\,,$ (8)
where $\mathds{1}$ is the indicator function. We obtain the final accuracy averaging over all batches in the dataset. For our evaluation we use $K=2$ and depict in Table 2 the results across all combinations of models and test datasets. While both “COCOLA MoisesDB” and “COCOLA Slakh2100” perform only slightly better than a random choice, “COCOLA CocoChorales” features improved performance. Finally, combining the three dataset, we obtain an accuracy of over 90% on all datasets, showcasing generalization with 90.43% on the held-out MUSDB18-HQ.
5.2 Accompaniment Generation Evaluation
For the accompaniment generation evaluation, we compare the MSDM model [11] with the proposed CompoNet (Section 3.3). We train CompoNet on MUSDB18-HQ and Slakh2100 (restricted to Bass, Drums, Guitar and Piano stems at test time). We also consider a Random baseline, where, for a given input, we output a random sub-mix from a different test track. We generate 200 chunks for both datasets and models, conditioning on random stem subsets of test tracks and querying a subset of the complementary. The chunks are $\sim$6s / 10.24s long on MUSDB18-HQ / Slakh2100. Given that MSDM tends to generate silence, we sample 12 candidate tracks for each generated track, selecting the one with the highest $L^{2}$ norm. We compare the COCOLA score in Eq. (4) with the FAD [16, 70] metric (interpreted as a sub-FAD [15, 11]) computed with CLAP [9], EnCodec [71], and VGGish [72] backbones.
We showcase the results in Table 3. With the FAD metrics, the model assigns the best score to the Random baseline. This behavior can be explained by considering that the Random outputs are real data, and the FAD evaluates well the perceptual quality. At the same time, it fails to assess the coherence between the tracks, and tends to score MSDM better. With the “COCOLA All” score, Random is the lowest, while CompoNet scores best. CompoNet has the highest value on “COCOLA CocoChorales” which, however, scores MSDM lower than Random. We remember that “COCOLA All” achieves a higher accuracy in Table 2, resulting in more reliability. On Ground Truth, we compute upper-bounds of the COCOLA score using real positive pairs. CompoNet better approaches these bounds.
6 Conclusion
In this paper we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems. At the same time, we have introduced CompoNet, a novel compositional model for music based on ControlNet that can solve a wide range of tasks simultaneously. Finally, we have evaluated CompoNet using COCOLA, proposing a new way of assessing accompaniment generation.
We plan to improve the quality of COCOLA by training on additional stem-level datasets [73] or using data obtained by pre-separating [15] larger realistic music datasets [74]. In future work, we would also like to explore inference-side methods that can guide the diffusion process using COCOLA as a likelihood function, offering an alternative (or additional) loss for the GMSDI method [46].
7 Acknowledgments
The authors were partially supported by the ERC grant no. 802554 (SPECGEO), PRIN 2020 project no. 2020TA3K9N (LEGO.AI), PRIN 2022 project no. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001), PNRR MUR project no. PE0000013-FAIR, and AWS Cloud Credit for Research program.

### v3

@@
 We present COCOLA (Coherence-Oriented Contrastive Learning for Audio), a contrastive learning method for musical audio representations that captures the harmonic and rhythmic coherence between samples.
-Our method operates at the level of stems (or their combinations) composing music tracks and allows the objective evaluation of compositional models for music in the task of accompaniment generation.
-We also introduce a new baseline for compositional music generation called CompoNet, based on ControlNet, generalizing the tasks of MSDM, and quantify it against the latter using COCOLA.
-We release all models trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
-1 Introduction
-Recently, there have been significant advances in music generation in the continuous domain [1, 2, 3, 4, 5], thanks to the impressive development of generative models [6, 7, 8].
+Our method operates at the level of the stems composing music tracks and can input features obtained via Harmonic-Percussive Separation (HPS).
+COCOLA allows the objective evaluation of generative models for music accompaniment generation, which are difficult to benchmark with established metrics.
+In this regard, we evaluate recent music accompaniment generation models, demonstrating the effectiveness of the proposed method.
+We release the model checkpoints trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
+Index Terms:
+Contrastive learning, generative AI, music information retrieval, performance evaluation
+I Introduction
+Recently, there have been significant advances in music generation in the continuous domain [1, 2, 3, 4], thanks to the impressive development of generative models [5, 6].
 In addition to producing high-quality tracks of increasing length [3], these models offer precise semantic control through textual conditioning [7, 8].
 However, they are limited as tools for musical composition, since they output a final mix containing all stems.
-To overcome this, a new range of compositional generative models is emerging [11, 12, 13], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive.
-The most important application of these models is accompaniment generation, where, given multiple conditioning sources (combined or not), the model is asked to output a new set (or a mixture) of coherent stems.
-Although previous models could generate accompaniments [14, 15], they could not be used iteratively (acting sequentially on the model’s outputs) in a composition process.
+To overcome this, a new range of compositional generative models for music is emerging [9, 10, 11], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive.
+The most important application of these models is accompaniment generation, where, given multiple conditioning sources, the model is asked to output a coherent stem that should combine with them in a harmonically and rhythmically consistent way, as opposed to being dissonant.
+While pioneering models [12, 13] could generate waveform accompaniments, their output is limited to a single stem category, and as such they cannot be used iteratively (acting sequentially on the model’s outputs) in a composition process.
 A significant problem with this line of research is the lack of an objective metric for quantifying the coherence of the generated outputs w.r.t.
@@
 However, this metric is not optimal for assessing coherence, as it focuses on global quality instead of the level of harmony and rhythm shared by constituent stems.
-Figure 1: Illustration of COCOLA score. COCOLA is a contrastive model able to estimate the coherence between instrumental tracks and generated accompaniments.
-To this end, we propose a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1).
+Figure 1: Illustration of COCOLA Score. COCOLA is a contrastive model that estimates the coherence between instrumental tracks and generated accompaniments.
+To this end, we propose and release [^1] https://github.com/gladia-research-group/cocola a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1).
 The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub-mixtures of stems) and minimizing it on sub-components belonging to different windows.
-With the model, we define a COCOLA score as the similarity between conditioning tracks and accompaniments in the embedding space.
-Additionally, given the scarcity of open-source compositional music models (to our knowledge, only MSDM is available publicly [11]), we introduce and release [^1] https://github.com/gladia-research-group/cocola an improved latent diffusion model called CompoNet, based on ControlNet [17], which generalizes the tasks of previous compositional models.
-We benchmark CompoNet against MSDM with the COCOLA score and FAD [16] on accompaniment generation, showcasing better performance.
-After discussing related work in Section 2, we introduce COCOLA and CompoNet in Section 3.
-We describe the experimental setup in Section 4 and present the results in Section 5.
-We conclude the article in Section 6.
+With the model, we define a COCOLA Score as the similarity between conditioning tracks and accompaniments in the embedding space, which we use to evaluate accompaniment generation models.
+After discussing related work in Section II, we introduce COCOLA in Section III.
+We describe the experimental setup in Section IV and present the results in Section V.
+We conclude the article in Section VI.
 Figure 2: The COCOLA training procedure (single stem case).
@@
 As a second step, we randomly select two distinct stems in each window.
-For example, in the first window we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums).
+For example, in the first window, we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums).
 Thus, we embed all selected stems with the COCOLA encoder $f_{\theta}$, obtaining latent representations.
@@
 Finally, we compute the contrastive loss (Eq.
-(3)) considering embedings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
-2 Related Work
-2.1 Contrastive Methods for Audio
+(3)) considering embeddings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
+II Related Work
+II-A Contrastive Methods for Audio
 Contrastive learning [15, 16] can be formulated both as a supervised or self-supervised problem.
-Supervised contrastive learning methods are typically cross-modal, requiring labeled information alongside audio data.
-In early works, the labeled information was in the form of simple tags, while the loss used to align embeddings of audio segments and tags was the triplet loss [20].
-Within the same data setting,[21] used the contrastive loss of SimCLR [22].
-With the advent of the transformer architecture [23], using complex sentences instead of simple tags became feasible.
-MuLaP [24] is the first model to train a common representation between audio and sentences in the musical domain.
-In such work, the audio and text are processed by a joint transformer encoder, conveying information about the two modalities through cross-attention layers.
-Although it is not a contrastive model per se, an audio-text matching loss uses negative examples to encourage the model to focus on aligned pairs.
-More recent works [25, 26, 9, 27], consider separate textual and audio encoders, which makes it possible to use the two branches independently at inference time.
-Self-supervised representation learning methods [28, 29, 30, 31] build embedding spaces targeting structural information extracted from the audio data itself.
-In [32], the authors build positive examples for a triplet loss by augmenting with Gaussian noise, time and frequency translations, and sampling with time proximity.
-They also consider example mixing.
-While we compare coherent mixes in our method (Section 3.1), in [32], positive pairs are not coherence-related (e.g., mixing siren and dog sounds).
-As in the supervised case, following [22], multi-class cross-entropy losses are employed [33, 34, 35].
-In COLA [33], the authors train an embedding model with contrastive loss using the simple criterion of sampling positive pairs only from the same audio track (still employing Gaussian noise), outperforming a fully supervised baseline in a plethora of tasks.
+Supervised contrastive learning methods often rely on cross-modal approaches, requiring labeled data alongside audio.
+Early works [17] employ the contrastive loss [18] for aligning audio with simple labels.
+MuLaP [19] was the first model to jointly represent audio and complex text using a joint transformer encoder with cross-attention layers.
+Recent models [20, 7, 21] have adopted separate text and audio encoders, allowing independent use of these branches during inference.
+Self-supervised representation learning methods [22, 23] create embeddings from structural information in audio data.
+In [24], positive examples for triplet loss are built using Gaussian noise, time/frequency translations, and time proximity, without ensuring coherence (e.g., mixing unrelated sounds).
+Following methods [25] adapted the contrastive loss.
+COLA [26] uses the simple criterion of sampling positive pairs only from the same audio track, outperforming a supervised baseline.
 [27] pairs mixtures with sources extracted via source separation.
-The proposed COCOLA method shares aspects of both supervised and self-supervised approaches.
-Given that stems are pre-separated, we cannot consider the method purely self-supervised.
-At the same time, we process such data with a uni-modal encoder, as is the case of self-supervised methods.
-2.2 Compositional Waveform Music Generation
-Compositional music generation in the waveform domain (as opposed to symbolic domains such as MIDI or sheet music [37]) was introduced by [11], proposing the Multi-Source Diffusion Model (MSDM).
-Such a model captures the joint distribution of a fixed set of sources (e.g., Bass, Drums, Guitar, and Piano) using a (score-based) diffusion model [38, 39, 7, 6, 40].
-At inference time, it is possible to perform unconditional generation of all sources, accompaniment generation of a subset given the complement, and source separation.
-Following this initial work, AUDIT [41] proposes a diffusion model conditioned by a T5 Encoder [10], trained with instructions that allow the addition, removal (drop), and replacement of sources in an input audio mixture.
-This model operates on general audio signals with weak dependencies between the sources (e.g., environmental sounds).
-While MSDM is an unconditional generative model that processes single sources in parallel, AUDIT is a conditional generative model that processes mixtures sequentially.
-InstructME [12] introduces AUDIT in the musical setting of MSDM, where sources are highly interdependent.
-Besides conditioning on chords, the fundamental difference lies in how the audio input is provided to the model: while in AUDIT, the input and output are two channels of the tensor that the diffusion model processes (the input is inpainted at inference time, similarly to how accompaniment generation is performed in MSDM), in InstructME, the input is processed by a convolutional network that follows the structure of the diffusion model U-Net [42] encoder.
-The features processed by the latter are aggregated with the features of the U-Net encoder (authors do not specify if they sum or inject the features).
-StemGen [13] generates a single accompaniment source, given an instrument tag and an input audio mixture, via a masked music language model [4].
-Another line of research focuses on inference-only methods for compositional music tasks, given pre-trained generative models.
-Based on generative source separation via Bayesian inference [43, 44, 45], Generalized Multi-Source Diffusion Inference (GMSDI) [46] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
-Our proposed CompoNet baseline (Section 3.3) is a sequential conditional model that, unlike previous models, can perform all compositional music generation tasks (Figure 3).
-Differently than InstructME, the model is conditioned via a ControlNet adapter [17], which enables fine-tuning of diffusion models pre-trained with a large quantity of mixtures.
-Table 1 summarizes a comparison between CompoNet and current music compositional models.
-3 Method
-3.1 Stem-Level Contrastive Learning
+The proposed COCOLA method combines elements of both supervised and self-supervised approaches.
+Since stems are pre-separated, it is not fully self-supervised, but like self-supervised methods, it processes data using a uni-modal encoder.
+II-B Waveform Music Accompaniment Generation Models
+MSDM [9], based on diffusion models [6, 5] was the first model able to generate waveform accompaniments belonging to different stem classes.
+Following, state-of-the-art models [11, 28] took the form of conditional generative models which respond with an output stem to an input track.
+Based on generative source separation via Bayesian inference [29, 30, 31], GMSDI [32] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
+Another category of models [10, 33], following AUDIT [34], perform music editing by modifying the input track.
+Such models can add instruments to a track, but given the generative nature of the model, isolating such stems via difference with the input or source separation can introduce artefacts.
+III Method
+III-A Stem-Level Contrastive Learning
 In our setting, we have access to a dataset $D=\{\bar{\mathbf{x}}^{k}\}_{k=1,\dots,\bar{K}}$ containing $\bar{K}$ musical tracks $\bar{\mathbf{x}}^{k}$, each separated into a variable number $N$ of individual stems $\bar{\mathbf{x}}^{k}_{n}$, i.e., $\bar{\mathbf{x}}^{k}=\{\bar{\mathbf{x}}^{k}_{n}\}_{n=1,\dots,N}$.
@@
 We define the sub-mixes $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ by summing the stems in $X_{1}^{k},X_{2}^{k}$:
-$\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}$ (1)
+$\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}\,.$ (1)
 When $X_{1}^{k},X_{2}^{k}$ are singletons, the sub-mixes are simply two stems in the window (single stem case).
-We work with sub-mixes because current compositional music generation methods [13] operate over them, including our proposed CompoNet (see Section 3.3).
 Like in COLA[26], we use a convolutional audio-only encoder [^2] In our notation, we incorporate into $f_{\theta}$ any domain transform preceding or following the convolutional network operations, like the (pre) mel-filterbank map and the (post) projection head $g$ in COLA.
@@
 We depict the training procedure of COCOLA in Figure 2 for the single stem case.
-Figure 3: Inter-stem compositional generation tasks. $Y$ and $X$ represent the input and output stem sub-sets, respectively.
-In the COLA training procedure, the positive pairs are (fully mixed) windows belonging to the same track.
-In COCOLA, they are sub-mixes belonging to the same window.
-As such, we allow for negative pairs belonging to the same track but in different windows.
-The $r$ ratio has to be chosen well to avoid strong overlaps between windows in the same track.
-In that case, we could potentially consider (nearly) coherent sub-mixes as negative pairs.
-Model Task Methodology Input Output Coherence
-UG AG SS EA ER
-MSDM [11] ✓(MS) ✓(MS) Training / Supervised Multi / Source Multi / Source
-GMSDI [46] ✓(MS) ✓(MS) Inference / Weakly Supervised Multi / Sub-mix Multi / Sub-mix
-StemGen [13] ✓(1S) Training / Supervised Single / Sub-mix Single / Source
-Audit [41] ✓(Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
-InstructME [12] ✓(Extract 1S; Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
+TABLE I: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
+Test Dataset
+Model MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
 […]
-Table 1: Compositional audio models comparison.
-The various tasks are illustrated in Figure 3.
-1S vs MS: the task operates on one vs multiple sources at a time.
-Multi vs Single on Input / Output: the model accepts multiple vs single inputs / outputs.
-Source vs Sub-mix on Input / Output: the model processes single sources or sub-mixes as inputs / outputs.
-3.2 COCOLA Score
-Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output).
-The model’s variables can be either a set of stems or sub-mixes.
-Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$.
-We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems).
-We define the COCOLA score (CCS) between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as:
-$\text{CCS}(\mathbf{y},\tilde{\mathbf{x}})=\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})\,,$ (4)
-the similarity (Eq. (2)) between their embeddings. The described procedure is depicted in Figure 1.
-3.3 CompoNet
-In order to showcase the utility of the COCOLA score as a metric for measuring the coherence of generated accompaniments, we propose an improved compositional model for music called CompoNet and compare it with the MSDM baseline in Section 5.
-CompoNet first pre-trains a latent diffusion model $\epsilon_{\phi}$, based on U-Net architecture [42] conditioned via cross-attention layers [23], on a large dataset of tuples $(\mathbf{m},\mathbf{c})$ comprising audio mixtures $\mathbf{m}$ and relative textual descriptions $\mathbf{c}$.
-The mixtures $\mathbf{m}$ are mapped to latent vectors $\mathbf{z}=E^{\text{VAE}}(\mathbf{m})$ using a pre-trained VAE encoder [47], while the text descriptions are mapped to a continuous sequence $\mathbf{s}=E^{\text{TXT}}(\mathbf{c})$ using a text encoder (e.g., [10]).
-Following the DDPM formulation [7], the model is trained to reverse the forward Gaussian noising process given by:
-$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\,,$ (5)
-where $t\in[0,T]$ is a time index (with $T$ the maximum time step), $\bar{\alpha}_{t}$ is defined by integrating a noise schedule [7], and $\boldsymbol{\epsilon}$ is sampled from the standard Gaussian distribution.
-The model is trained with denoising score matching [39]:
-$\min_{\phi}\mathbb{E}_{\mathbf{z},\boldsymbol{\epsilon},\mathbf{s},t}\left[\|\boldsymbol{\epsilon}-\epsilon_{\phi}(\mathbf{z}_{t},\mathbf{s},t)\|^{2}_{2}\right]\,,$ (6)
-with $\mathbf{z}_{t}$ obtained via Eq. (5).
-In a second phase, we fine-tune a ControlNet [17] adapter $C_{\psi}$ for tackling all compositional musical tasks (Figure 3) with a single model.
-The U-Net $\epsilon_{\phi}$ comprises an encoder, bottleneck, decoder structure $\epsilon_{\phi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E}}$.
-The ControlNet adapter is defined as $C_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E_{\phi_{E}}(\mathbf{z}_{t}+\text{conv}_{\text{in}}(\mathbf{w}),\mathbf{s},t)$, where $\text{conv}_{\text{in}}$ is a zero-initialized convolutional layer, and $\mathbf{w}$ is a latent (VAE) embedding of an external audio input.
-The ControlNet adapter outputs the set of processed features $\{C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})\}_{i=1,\dots,I}$ for each layer of $E_{\phi_{E}}$, with $I$ the total number of layers.
-The full ControlNet conditional architecture is defined as $\epsilon_{\phi,\psi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E},\psi}$ with $E_{\phi_{E},\psi}$ combining the encoder and ControlNet adapter features at each layer $i$ with zero-initialized convolutional layers $\text{conv}_{i}$:
-$E^{i}_{\phi_{E},\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E^{i}_{\phi_{E}}(\mathbf{z}_{t},\mathbf{s},t)+\text{conv}_{i}(C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w}))\,.$
-While we have described the general ControlNet architecture, we still have to describe how we train it in CompoNet, namely, the roles of the $\mathbf{z},\mathbf{w}$ and $\mathbf{s}$ variables.
-Iterating over a dataset containing tuples $(\mathbf{x},\mathbf{t})$ with multi-stem tracks $\mathbf{x}=\{\mathbf{x}_{n}\}_{n=1,\dots,N}$ and tag descriptions $\mathbf{t}=\{\mathbf{t}_{n}\}_{n=1,\dots,N}$ for each stem, we sample from $\mathbf{x}$ two arbitrary subsets of stems $Y,X\subseteq\mathbf{x}$, with $|X|>0$.
-$Y$ contains input stems while $X$ contains output stems.
-The topological relationships between such subsets define all possible compositional tasks, as depicted in Figure 3.
-While previous models partially solve some tasks (see Table 1), ours is the first to solve all of them simultaneously.
-We proceed like in Eq.
-(1) and mix the sources in $Y$ and $X$, obtaining $\mathbf{m}_{Y}$ and $\mathbf{m}_{X}$, respectively.
-Afterward, we encode them in the VAE latent space, defining $\mathbf{z}=E^{\text{VAE}}(\mathbf{m}_{X})$ and $\mathbf{w}=E^{\text{VAE}}(\mathbf{m}_{Y})$.
-We define the following prompt $\mathbf{s}$:
-$\mathbf{s}=E^{\text{TXT}}(\mathbf{t}_{Y_{1}},\dots,\mathbf{t}_{Y_{|Y|}},\verb|SEP|,\ \mathbf{t}_{X_{1}},\dots,\mathbf{t}_{X_{|X|}})\,,$ (7)
-specifying input and output mixture tags separated by a special token SEP.
-Having specified the required inputs and outputs, we train $\epsilon_{\phi,\psi}$ via Eq.
-(6), optimizing only the $\psi$ weights.
-The $\mathbf{s}$ prompt instructs the model which task to perform based on the specified stem tags.
-The user knows such information during the composition process in a generative DAW [48].
-While we only test CompoNet on accompaniment generation against MSDM using COCOLA, we want the models to share similar abilities for a fair comparison.
-4 Experimental Setup
-Test Dataset
-Train Dataset MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
 MoisesDB [35] 52.56 53.01 51.22 60.32
@@
 All 90.43 93.06 90.06 99.89
-Table 2: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
-4.1 Datasets
-In our experiments, we use four different stem-separated public datasets for training COCOLA and fine-tuning CompoNet.
-The datasets are MUSDB18-HQ [52], MoisesDB [49], Slakh2100 [50] and CocoChorales [51].
-MUSDB18-HQ [52] is the uncompressed version (in WAV format) of the MUSDB18 dataset, initially introduced in [53].
-This dataset is a standard for evaluating music source separation systems.
-It comprises 150 tracks – 100 for training and 50 for testing – totaling approximately 10 hours of professional-quality audio.
-Each track is divided into four stems: Bass, Drums, Vocals, and Other, with “Other” covering any components not classified under the first three categories.
-MoisesDB [49] features 240 music tracks across diverse genres and artists, accumulating more than 14 hours of music.
-Unlike MUSDB18-HQ, MoisesDB is a genuine multi-track dataset, offering a two-tier taxonomy of 11 distinct stems.
-Each stem in this dataset includes more detailed type annotations (e.g., Guitar might be labeled as Acoustic Guitar or Electric Guitar).
-Not having pre-computed splits, we set a custom 0.8 (train) / 0.1 (validation) / 0.1 (test) split.
-Slakh2100 [50] is synthesized from the Lakh MIDI Dataset v0.1 [54] employing high-quality sample-based virtual instruments.
-It features 2100 tracks organized into 1500 tracks for training, 375 for validation, and 225 for testing, together amounting to 145h of audio.
-The tracks are annotated into 34 stem categories.
-While such a dataset contains an order of magnitude more data than MUSDB18-HQ and MoisesDB, it does not share the same level of realism as the latter, being the tracks synthesized from MIDI.
-CocoChorales [51] is a chorale audio music dataset created through a synthesis process like Slakh2100.
-However, it comprises a substantially vaster collection of 240000 tracks, extending over 1411 hours of audio data.
-It is produced by generating symbolic notes via a Coconet model, performing their synthesis with MIDI-DDSP [55].
-This dataset is richly annotated, featuring details on performance attributes and synthesis parameters.
-CocoChorales includes a diverse range of 13 instruments spanning Strings, Brass, Woodwind, and Random ensembles.
-In our experiments we use the tiny version, comprising 4000 tracks.
-4.2 Model Implementation
-To implement the COCOLA encoder $f_{\theta}$, we follow [33] and employ the EfficientNet-B0 [56] convolutional architecture followed by a linear projection layer, operating on the mel-filterbank audio representation.
-The embedding dimension is 512.
-With respect to the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
-For CompoNet, we employ the AudioLDM2 [57, 58] architecture.
-Since the authors pre-train the model on a large array of datasets [59, 60, 61, 62, 63, 64, 65, 66], we skip the pre-training phase in Section 3.2 and directly fine-tune a ControlNet adapter based on the AudioLDM 2-Large checkpoint [^3] https://huggingface.co/cvssp/audioldm2-large.
-During fine-tuning, we directly pass the conditioning prompt in Eq.
-7 to the text-embedding mechanism of AudioLDM2 (based on CLAP [9, 27], T5 [10], and GPT2 [67]), conditioning both the U-Net and the ControlNet adapter.
-100100100100200200200200300300300300400400400400500500500500600600600600700700700700800800800800000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Training stepTrain accuracyCosine similarityBilinear similarity
-Figure 4: Cosine vs bilinear similarity during training.
-Training is performed with “COCOLA All”.
-The accuracy metric is defined in Eq.
-(8).
-We adopt the bilinear similarity in Eq.
-(2), given its improved performance.
+w\ HPS
+All 93.87 93.67 94.27 99.68
+TABLE II: Classification accuracy tests (%) with COCOLA HPS All varying the values of $K$.
+Test Dataset
+K MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
+8 78.02 73.68 79.72 98.27
+16 70.32 62.17 72.62 96.67
+64 54.33 34.04 59.35 90.67
+TABLE III: Comparison between music accompaniment generation models.
 Method
@@
 VGGish
+[…]
+Percussive + Harmonic
 COCOLA Score $\uparrow$
-All
+Percussive
 COCOLA Score $\uparrow$
-CocoChorales
-Slakh2100
-MSDM [11] 0.23 92.81 2.01 3.31 0.72
-CompoNet 0.30 106.23 3.20 13.50 8.32
-Random 0.064 51.44 0.16 0.069 1.06
+Harmonic
+MoisesDB Test
+Random 0.072 19.37 0.38 50.53 52.34 53.86
 […]
-MUSDB18-HQ
-MSDM [11] 0.29 148.09 2.36 11.61 3.37
-CompoNet 0.37 130.04 2.14 11.94 7.21
-Random 0.11 100.25 0.35 4.40 3.47
+SA ControlNet [4, 42] 0.15 170.21 2.59 55.11 55.89 57.34
+Diff-A-Riff [28] 0.14 12.16 0.90 57.34 58.00 60.00
 […]
-Table 3: Comparison between MSDM and CompoNet.
-4.3 Training Details
+MUSDB18-HQ Test
+Random 0.087 6.04 0.49 49.49 51.27 51.32
+[…]
+SA ControlNet [4, 42] 0.12 12.14 0.66 56.41 57.48 58.82
+Diff-A-Riff [28] 0.20 119.41 1.68 53.79 54.75 56.08
+[…]
+In the COLA training procedure, the positive pairs are (fully mixed) windows belonging to the same track.
+In COCOLA, they are sub-mixes belonging to the same window.
+As such, we allow for negative pairs belonging to the same track but in different windows.
+The $r$ ratio has to be chosen well to avoid strong overlaps between windows in the same track.
+In that case, we could potentially consider (nearly) coherent sub-mixes as negative pairs.
+In addition to the standard mel-filterbank representation of the input (as defined in COLA), we propose a variant where the COCOLA encoder operates on a pair of mel spectrogram features obtained via harmonic-percussive separation (HPS) [38].
+As shown in Section V, this factorized input not only improves performance but also lets us better interpret which factor (harmony or rhythm) is more coherent.
+III-B COCOLA Score
+Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output).
+The model’s variables can be either a set of stems or sub-mixes.
+Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$.
+We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems).
+We define the COCOLA Score between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as $\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})$, the similarity (Eq.
+(2)) between their embeddings.
+The described procedure is depicted in Figure 1.
+Figure 3: Correlation plot between subjective scores (MOS) ($x$-axis) and COCOLA Scores ($y$-axis).
+IV Experimental Setup
+IV-A Datasets
+In our experiments, we use four different stem-separated public datasets for training COCOLA.
+The datasets are MUSDB18-HQ [39], MoisesDB [35], Slakh2100 [36] and CocoChorales [37].
+For MoisesDB, not having pre-determined splits, we set custom 0.8 (train) / 0.1 (validation) / 0.1 (test) splits.
+For CocoChorales, we use the tiny version, comprising a subset of 4000 tracks.
+IV-B Model Implementation
+To implement the COCOLA encoder $f_{\theta}$, we follow the COLA framework [26] and employ the EfficientNet-B0 [40] (using two input channels for HPS) convolutional architecture followed by a linear projection layer.
+The embedding dimension is 512.
+Differently from the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
+IV-C Training Details
 All COCOLA models are trained on an NVIDIA RTX 4070 Super with 12GB of VRAM.
 Each training batch contains 32 audio chunks of 5s (16kHz).
-We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer [68] with a $10^{-3}$ learning rate.
-We add Gaussian noise to positive samples as a data augmentation method, with $\sigma=10^{-3}$.
-We tried training with cosine similarity [22] but we reported (Figure 4) lower performance, corroborating [33].
-CompoNet models are also fine-tuned on NVIDIA GPUs (RTX 4080 16GB on MUSDB18-HQ; A10G, 24GB on Slakh2100).
-The training batches have size 2 (MUSDB18-HQ) and 5 (Slakh2100) and contain windows of 10.24s (16kHz).
-We train with Adam [68] with a $10^{-4}$ learning rate.
-We do not use the empty prompt for classifier-free guidance [69] during fine-tuning.
-5 Experiments
-We employ four COCOLA encoder models in our experiments : “COCOLA MoisesDB”, “COCOLA Slakh2100”, “COCOLA CocoChorales” and “COCOLA All”.
+We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer with a $10^{-3}$ learning rate.
+As a data augmentation method, we add Gaussian noise to positive samples, with $\sigma=10^{-3}$.
+During the training with HPS, masking is randomly applied to either one of the two channels or none. This approach allows the model to learn representations from either component independently.
+V Experiments
+In our experiments, we trained four COCOLA encoder models without HPS: “COCOLA MoisesDB”, “COCOLA Slakh2100”, “COCOLA CocoChorales” and “COCOLA All”.
 The first three are trained on the homonym datasets, while the last one is trained on all three combined.
-For the “COCOLA CocoChorales” we use all ensables while on “COCOLA All” we use only the Random ensamble for a more balanced partitioning, with respect to the other datasets.
+For the “COCOLA CocoChorales” we use all ensembles, while on “COCOLA All” we use only the Random ensemble for a more balanced partitioning with respect to the other datasets.
+Additionally, we train a “COCOLA HPS All” model employing HPS with the same data as “COCOLA ALL”.
 MUSDB18-HQ, being the smallest dataset, is used as a held-out test dataset for studying generalization.
-5.1 Coherent Sub-Mix Classification
-We cross-test the performance of all COCOLA models, performing classification of coherent pairs on the test split of our datasets.
+V-A Coherent Sub-Mix Classification
+We cross-test the performance of COCOLA models trained without HPS by classifying coherent pairs on the test splits of our datasets.
 More specifically, given an encoder $f_{\theta}$, we iterate a test set, collecting at each step a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$.
-Following the steps in Section 3.1 we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$.
+Following the steps in Section III-A we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$.
 We define the accuracy over a batch as:
@@
 where $\mathds{1}$ is the indicator function.
-We obtain the final accuracy averaging over all batches in the dataset.
-For our evaluation we use $K=2$ and depict in Table 2 the results across all combinations of models and test datasets.
-While both “COCOLA MoisesDB” and “COCOLA Slakh2100” perform only slightly better than a random choice, “COCOLA CocoChorales” features improved performance.
-Finally, combining the three dataset, we obtain an accuracy of over 90% on all datasets, showcasing generalization with 90.43% on the held-out MUSDB18-HQ.
-5.2 Accompaniment Generation Evaluation
-For the accompaniment generation evaluation, we compare the MSDM model [11] with the proposed CompoNet (Section 3.3).
-We train CompoNet on MUSDB18-HQ and Slakh2100 (restricted to Bass, Drums, Guitar and Piano stems at test time).
-We also consider a Random baseline, where, for a given input, we output a random sub-mix from a different test track.
-We generate 200 chunks for both datasets and models, conditioning on random stem subsets of test tracks and querying a subset of the complementary.
-The chunks are $\sim$6s / 10.24s long on MUSDB18-HQ / Slakh2100.
-Given that MSDM tends to generate silence, we sample 12 candidate tracks for each generated track, selecting the one with the highest $L^{2}$ norm.
-We compare the COCOLA score in Eq.
-(4) with the FAD [16, 70] metric (interpreted as a sub-FAD [15, 11]) computed with CLAP [9], EnCodec [71], and VGGish [72] backbones.
-We showcase the results in Table 3.
-With the FAD metrics, the model assigns the best score to the Random baseline.
-This behavior can be explained by considering that the Random outputs are real data, and the FAD evaluates well the perceptual quality.
-At the same time, it fails to assess the coherence between the tracks, and tends to score MSDM better.
-With the “COCOLA All” score, Random is the lowest, while CompoNet scores best.
-CompoNet has the highest value on “COCOLA CocoChorales” which, however, scores MSDM lower than Random.
-We remember that “COCOLA All” achieves a higher accuracy in Table 2, resulting in more reliability.
-On Ground Truth, we compute upper-bounds of the COCOLA score using real positive pairs.
-CompoNet better approaches these bounds.
-6 Conclusion
-In this paper we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems.
-At the same time, we have introduced CompoNet, a novel compositional model for music based on ControlNet that can solve a wide range of tasks simultaneously.
-Finally, we have evaluated CompoNet using COCOLA, proposing a new way of assessing accompaniment generation.
-We plan to improve the quality of COCOLA by training on additional stem-level datasets [73] or using data obtained by pre-separating [15] larger realistic music datasets [74].
-In future work, we would also like to explore inference-side methods that can guide the diffusion process using COCOLA as a likelihood function, offering an alternative (or additional) loss for the GMSDI method [46].
-7 Acknowledgments
-The authors were partially supported by the ERC grant no.
-802554 (SPECGEO), PRIN 2020 project no.
-2020TA3K9N (LEGO.AI), PRIN 2022 project no.
-2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001), PNRR MUR project no.
-PE0000013-FAIR, and AWS Cloud Credit for Research program.
+We compute the final accuracy by averaging across all batches.
+Using $K=2$, Table I shows results for various model and test dataset combinations.
+While ”COCOLA MoisesDB” and ”COCOLA Slakh2100” perform slightly better than random, ”COCOLA Coco-Chorales” shows improved performance.
+Combining all three datasets achieves over 90% accuracy.
+Adding HPS to “COCOLA All”, all metrics improve except on CocoChorales (which is already high), reaching 93.87% on the held-out MUSDB18-HQ and showcasing the usefulness of the factorized representation.
+We also investigate the performance of “COCOLA HPS All” increasing $K$ in Table II.
+V-B Accompaniment Generation Evaluation
+For the music accompaniment generation evaluation, we compare GMSDI [32], Diff-A-Riff [28], and Stable Audio ControlNet (SA ControlNet).
+The latter is an implementation of a DiT-based [41] ControlNet [42] for Stable Audio Open [4], which we release [^3] https://github.com/EmilianPostolache/stable-audio-controlnet, given the lack of state-of-the-art open-source models for music accompaniment generation (only MSDM [9] is available, but it is trained on Slakh2100 [36] not being applicable in realistic settings).
+GMSDI is trained on MTG-Jamendo [43], Diff-A-Riff on a proprietary dataset and we finetune SA ControlNet on MoisesDB.
+We also consider a Random baseline, where, for a given input, we output a random stem from a different test chunk.
+We evaluate the models both on the MoisesDB and MUSDB18-HQ test splits.
+We sample 168 chunks ($\sim$47.55s) from the MoisesDB test split and 236 chunks of the same length from the MUSDB18-HQ test split, using random stem subsets as input and querying complementary stems, excluding vocals from the output and the ”Other” class in MUSDB18-HQ.
+SA ControlNet processes the full chunk, while GMSDI and Diff-A-Riff use shorter initial sub-chunks ($\sim$23.78s and $\sim$9.58s, respectively) due to reduced context windows.
+Outputs are cut to the initial $\sim$9.58s.
+We compare the COCOLA Score (based on ”COCOLA HPS All”) with the FAD [14, 44] metric (interpreted as a sub-FAD [13, 9]), computed using CLAP [7], EnCodec [45], and VGGish.
+Results in Table III show that FAD tends to favor the Random baseline, as it evaluates perceptual quality but not track coherence, which is better captured by COCOLA.
+On Ground Truth, we compute COCOLA upper bounds with real positive pairs.
+Diff-A-Riff performs best on MoisesDB, and SA ControlNet on MUSDB18-HQ.
+V-C MOS vs COCOLA Score
+To evaluate COCOLA’s effectiveness, we conducted a subjective test with 41 participants who rated (0-4) the coherence of (pitch) Shifted and (time) Stretched chunks.
+We also used Random, Ground Truth, Diff-A-Riff and SA ControlNet tracks from the previous experiment.
+We average the ratings (MOS) for each chunk.
+Figure 3 shows the correlation plot between MOS and COCOLA Score (using ”COCOLA HPS All”).
+The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$ indicating a good correlation with high statistical signficance.
+VI Conclusion
+In this paper, we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems.
+Both classification experiments and human scores demonstrate the efficacy of the method.
+We used COCOLA for benchmarking music accompaniment models, proposing a new evaluation metric for the task.
+We plan to improve the quality of COCOLA by training on additional stem-level datasets [46] or on pre-separated [13] large scale datasets [43].

### v4

@@
 We present COCOLA (Coherence-Oriented Contrastive Learning for Audio), a contrastive learning method for musical audio representations that captures the harmonic and rhythmic coherence between samples.
-Our method operates at the level of the stems composing music tracks and can input features obtained via Harmonic-Percussive Separation (HPS).
-COCOLA allows the objective evaluation of generative models for music accompaniment generation, which are difficult to benchmark with established metrics.
-In this regard, we evaluate recent music accompaniment generation models, demonstrating the effectiveness of the proposed method.
+Our method operates at the level of the individual stems composing music tracks and can input features obtained via Harmonic-Percussive Separation (HPS).
+COCOLA allows an objective evaluation of generative models for music accompaniment generation, which are difficult to benchmark with established metrics.
+In this regard, we evaluate recent music accompaniment generation models, demonstrating the effectiveness of our proposed method.
 We release the model checkpoints trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
@@
 To overcome this, a new range of compositional generative models for music is emerging [9, 10, 11], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive.
-The most important application of these models is accompaniment generation, where, given multiple conditioning sources, the model is asked to output a coherent stem that should combine with them in a harmonically and rhythmically consistent way, as opposed to being dissonant.
-While pioneering models [12, 13] could generate waveform accompaniments, their output is limited to a single stem category, and as such they cannot be used iteratively (acting sequentially on the model’s outputs) in a composition process.
-A significant problem with this line of research is the lack of an objective metric for quantifying the coherence of the generated outputs w.r.t.
+The main application of these models is accompaniment generation, where the model is required to produce a coherent stem that blends harmonically and rhythmically with multiple conditioning sources.
+While pioneering models [12, 13] can generate waveform accompaniments, their output is limited to a single stem category, and as such they cannot be used iteratively (acting sequentially on the model’s outputs) in a composition process.
+A significant problem with this line of research is the lack of objective metrics for quantifying the coherence of the generated outputs w.r.t.
 the inputs.
 For example, [9] proposes the sub-FAD metric as a multi-stem generalization of the FAD [14] protocol proposed in [13].
-However, this metric is not optimal for assessing coherence, as it focuses on global quality instead of the level of harmony and rhythm shared by constituent stems.
+However, this metric is suboptimal for assessing coherence, as it focuses on global quality instead of the level of harmony and rhythm shared by constituent stems.
 Figure 1: Illustration of COCOLA Score. COCOLA is a contrastive model that estimates the coherence between instrumental tracks and generated accompaniments.
 To this end, we propose and release [^1] https://github.com/gladia-research-group/cocola a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1).
-The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub-mixtures of stems) and minimizing it on sub-components belonging to different windows.
-With the model, we define a COCOLA Score as the similarity between conditioning tracks and accompaniments in the embedding space, which we use to evaluate accompaniment generation models.
+The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub mixtures of stems) and minimizing it on sub-components belonging to different windows.
+With the model, we define a COCOLA Score as the similarity between conditioning tracks and accompaniments in the embedding space, and use it to benchmark accompaniment generation models.
 After discussing related work in Section II, we introduce COCOLA in Section III.
@@
 Figure 2: The COCOLA training procedure (single stem case).
-We first randomly crop windows of size $L$ from a batch of $K$ tracks (depicted on the left).
-As a second step, we randomly select two distinct stems in each window.
-For example, in the first window, we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums).
-Thus, we embed all selected stems with the COCOLA encoder $f_{\theta}$, obtaining latent representations.
-For example, we obtain $\mathbf{h}^{1}_{1}$ and $\mathbf{h}^{1}_{2}$ from the first window.
-Finally, we compute the contrastive loss (Eq.
-(3)) considering embeddings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
+Windows of size $L$ are randomly cropped from $K$ tracks (left).
+Two distinct stems per window are randomly selected (e.g., Guitar $\mathbf{x}^{1}_{1}$ and Drums $\mathbf{x}^{1}_{3}$ in the first window), embedded using the COCOLA encoder $f_{\theta}$, yielding latent representations (e.g., $\mathbf{h}^{1}_{1}$ and $\mathbf{h}^{1}_{2}$).
+Contrastive loss (Eq.
+(3)) is computed with positive pairs within windows and negatives across windows.
 II Related Work
-II-A Contrastive Methods for Audio
-Contrastive learning [15, 16] can be formulated both as a supervised or self-supervised problem.
-Supervised contrastive learning methods often rely on cross-modal approaches, requiring labeled data alongside audio.
+II-A Contrastive Models for Audio
+Contrastive learning [15, 16] can be both formulated as a supervised or self-supervised problem.
+Supervised contrastive learning methods often rely on cross-modal approaches, requiring labeled data alongside the audio.
 Early works [17] employ the contrastive loss [18] for aligning audio with simple labels.
@@
 In [24], positive examples for triplet loss are built using Gaussian noise, time/frequency translations, and time proximity, without ensuring coherence (e.g., mixing unrelated sounds).
-Following methods [25] adapted the contrastive loss.
+Subsequent methods [25] adapted the contrastive loss.
 COLA [26] uses the simple criterion of sampling positive pairs only from the same audio track, outperforming a supervised baseline.
 [27] pairs mixtures with sources extracted via source separation.
-The proposed COCOLA method combines elements of both supervised and self-supervised approaches.
-Since stems are pre-separated, it is not fully self-supervised, but like self-supervised methods, it processes data using a uni-modal encoder.
+A line of research [28, 29, 30, 31] focuses on estimating the compatibility of loops for retrieval tasks.
+Although these approaches, like ours, aim to evaluate the coherence between audio tracks, they address a distinct task (retrieval) rather than generative model evaluation.
+Moreover, they have limitations: some of them do not extract individual stems [28], while others rely on fixed stem choices [29, 30, 31].
+In contrast, our method supports any combination of stems, offering greater flexibility.
 II-B Waveform Music Accompaniment Generation Models
-MSDM [9], based on diffusion models [6, 5] was the first model able to generate waveform accompaniments belonging to different stem classes.
+MSDM [9], based on diffusion models [6, 5] was the first model to generate waveform accompaniments belonging to different stem classes.
 Following, state-of-the-art models [11, 32] took the form of conditional generative models which respond with an output stem to an input track.
-Based on generative source separation via Bayesian inference [29, 30, 31], GMSDI [32] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
+Based on generative source separation via Bayesian inference [33, 34, 35], GMSDI [36] performs the same tasks as MSDM by enabling source separation during generation, requiring models trained solely on mixtures and text.
 Another category of models [10, 37], following AUDIT [38], perform music editing by modifying the input track.
-Such models can add instruments to a track, but given the generative nature of the model, isolating such stems via difference with the input or source separation can introduce artefacts.
+Such models can add instruments to a track, but given the generative nature of the model, isolating such stems via the difference with the input or source separation can introduce artifacts.
 III Method
@@
 Afterward, we select, for each $k$, two disjoint non-empty stem subsets $X_{1}^{k},X_{2}^{k}$ of $\mathbf{x}^{k}$.
-We define the sub-mixes $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ by summing the stems in $X_{1}^{k},X_{2}^{k}$:
+We define the submixes $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ as the sum of the stems in $X_{1}^{k},X_{2}^{k}$:
 $\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}\,.$ (1)
-When $X_{1}^{k},X_{2}^{k}$ are singletons, the sub-mixes are simply two stems in the window (single stem case).
+When $X_{1}^{k},X_{2}^{k}$ are singletons, the submixes are simply two stems in the window (single stem case).
 Like in COLA[26], we use a convolutional audio-only encoder [^2] In our notation, we incorporate into $f_{\theta}$ any domain transform preceding or following the convolutional network operations, like the (pre) mel-filterbank map and the (post) projection head $g$ in COLA.
 $f_{\theta}:\mathbb{R}^{L}\to\mathbb{R}^{d}$, mapping $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ to lower-dimensional embedding vectors $\mathbf{h}^{k}_{1}=f_{\theta}(\mathbf{m}^{k}_{1})$ and $\mathbf{h}^{k}_{2}=f_{\theta}(\mathbf{m}^{k}_{2})$, with $d$ the embedding dimension.
-The COCOLA training procedure maximizes the agreement between pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}$ of sub-mixes embeddings in the same window.
-It decreases it for pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}$ ($j\neq k$) of sub-mixes embeddings in different windows.
+The COCOLA training procedure maximizes the agreement between pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}$ of embeddings belonging to the same window while minimizing it for pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}$ ($j\neq k$) belonging to different windows.
 As in COLA, we use a bilinear similarity metric:
@@
 $\mathcal{L}=-\sum^{K}_{k=1}\log\frac{\exp(\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}))}{\sum^{K}_{j=1}\exp(\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}))}\,.$ (3)
-We depict the training procedure of COCOLA in Figure 2 for the single stem case.
 TABLE I: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
@@
 […]
-In the COLA training procedure, the positive pairs are (fully mixed) windows belonging to the same track.
-In COCOLA, they are sub-mixes belonging to the same window.
-As such, we allow for negative pairs belonging to the same track but in different windows.
-The $r$ ratio has to be chosen well to avoid strong overlaps between windows in the same track.
-In that case, we could potentially consider (nearly) coherent sub-mixes as negative pairs.
+In the COLA training procedure, positive pairs are (fully mixed) windows belonging to the same track.
+In COCOLA (Figure 2), they are submixes belonging to the same window.
+As such, we allow negative pairs to belong to the same track but in different windows.
+The $r$ ratio has to be chosen well to avoid strong overlaps between windows on the same track.
+In that case, we could potentially consider (nearly) coherent submixes as negative pairs.
 In addition to the standard mel-filterbank representation of the input (as defined in COLA), we propose a variant where the COCOLA encoder operates on a pair of mel spectrogram features obtained via harmonic-percussive separation (HPS) [43].
-As shown in Section V, this factorized input not only improves performance but also lets us better interpret which factor (harmony or rhythm) is more coherent.
+As shown in Section V, this factorized input improves performance and helps us better understand which factor (harmony or rhythm) exhibits greater coherence.
 III-B COCOLA Score
-Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output).
-The model’s variables can be either a set of stems or sub-mixes.
-Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$.
-We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems).
-We define the COCOLA Score between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as $\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})$, the similarity (Eq.
-(2)) between their embeddings.
-The described procedure is depicted in Figure 1.
-Figure 3: Correlation plot between subjective scores (MOS) ($x$-axis) and COCOLA Scores ($y$-axis).
+Equipped with the encoder $f_{\theta}$, we measure the coherence of accompaniments generated by a conditional model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable and $\mathbf{x}$ a set of stems or a submix.
+Given $\mathbf{y}$, the model outputs $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$.
+We embed both $\mathbf{y}$ and $\tilde{\mathbf{x}}$ via $f_{\theta}$, optionally summing stems, obtaining $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$.
+The COCOLA Score is then defined as $\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})$ (Eq.
+(2)), representing the similarity between the embeddings, as illustrated in Figure 1.
 IV Experimental Setup
 IV-A Datasets
-In our experiments, we use four different stem-separated public datasets for training COCOLA.
+In our experiments, we use four different stem-separated public datasets to train COCOLA.
 The datasets are MUSDB18-HQ [44], MoisesDB [39], Slakh2100 [40] and CocoChorales [41].
-For MoisesDB, not having pre-determined splits, we set custom 0.8 (train) / 0.1 (validation) / 0.1 (test) splits.
+For MoisesDB, which does not have predetermined splits, we set a custom 0.8 (train) / 0.1 (validation) / 0.1 (test) split.
 For CocoChorales, we use the tiny version, comprising a subset of 4000 tracks.
-IV-B Model Implementation
+Figure 3: Correlation plot between subjective scores (MOS) ($x$-axis) and COCOLA Scores ($y$-axis).
+IV-B Model and Training Details
 To implement the COCOLA encoder $f_{\theta}$, we follow the COLA framework [26] and employ the EfficientNet-B0 [45] (using two input channels for HPS) convolutional architecture followed by a linear projection layer.
@@
 Differently from the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
-IV-C Training Details
-All COCOLA models are trained on an NVIDIA RTX 4070 Super with 12GB of VRAM.
-Each training batch contains 32 audio chunks of 5s (16kHz).
+Each training batch contains 32 5s audio chunks (16kHz).
 We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer with a $10^{-3}$ learning rate.
 As a data augmentation method, we add Gaussian noise to positive samples, with $\sigma=10^{-3}$.
-During the training with HPS, masking is randomly applied to either one of the two channels or none. This approach allows the model to learn representations from either component independently.
+During HPS training, masking is applied randomly to one or both of the two channels.
+This approach allows the model to learn the representations of either component independently.
 V Experiments
@@
 The first three are trained on the homonym datasets, while the last one is trained on all three combined.
-For the “COCOLA CocoChorales” we use all ensembles, while on “COCOLA All” we use only the Random ensemble for a more balanced partitioning with respect to the other datasets.
-Additionally, we train a “COCOLA HPS All” model employing HPS with the same data as “COCOLA ALL”.
-MUSDB18-HQ, being the smallest dataset, is used as a held-out test dataset for studying generalization.
+For the “COCOLA CocoChorales” we use all the ensembles, while on “COCOLA All” we use only the Random ensemble for a more balanced partitioning with respect to the other datasets.
+Furthermore, we train a “COCOLA HPS All” using HPS with the same dataset as “COCOLA ALL”.
+To assess generalization, the smallest dataset, MUSDB18-HQ, is reserved as a held-out test set.
 V-A Coherent Sub-Mix Classification
 We cross-test the performance of COCOLA models trained without HPS by classifying coherent pairs on the test splits of our datasets.
-More specifically, given an encoder $f_{\theta}$, we iterate a test set, collecting at each step a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$.
+More specifically, given an encoder $f_{\theta}$, we iterate over a test set, collecting a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$ at each step.
 Following the steps in Section III-A we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$.
@@
 We compute the final accuracy by averaging across all batches.
-Using $K=2$, Table I shows results for various model and test dataset combinations.
+Using $K=2$, Table I shows the results for various combinations of models and test data sets.
 While ”COCOLA MoisesDB” and ”COCOLA Slakh2100” perform slightly better than random, ”COCOLA Coco-Chorales” shows improved performance.
@@
 V-B Accompaniment Generation Evaluation
-For the music accompaniment generation evaluation, we compare GMSDI [32], Diff-A-Riff [28], and Stable Audio ControlNet (SA ControlNet).
+For the evaluation of generated music accompaniments, we compare GMSDI [36], Diff-A-Riff [32], and Stable Audio ControlNet (SA ControlNet).
 The latter is an implementation of a DiT-based [46] ControlNet [42] for Stable Audio Open [4], which we release [^3] https://github.com/EmilianPostolache/stable-audio-controlnet, given the lack of state-of-the-art open-source models for music accompaniment generation (only MSDM [9] is available, but it is trained on Slakh2100 [40] not being applicable in realistic settings).
@@
 We also consider a Random baseline, where, for a given input, we output a random stem from a different test chunk.
-We evaluate the models both on the MoisesDB and MUSDB18-HQ test splits.
+We evaluate both models on the MoisesDB and MUSDB18-HQ test splits.
 We sample 168 chunks ($\sim$47.55s) from the MoisesDB test split and 236 chunks of the same length from the MUSDB18-HQ test split, using random stem subsets as input and querying complementary stems, excluding vocals from the output and the ”Other” class in MUSDB18-HQ.
 SA ControlNet processes the full chunk, while GMSDI and Diff-A-Riff use shorter initial sub-chunks ($\sim$23.78s and $\sim$9.58s, respectively) due to reduced context windows.
-Outputs are cut to the initial $\sim$9.58s.
+The outputs are cut to the initial $\sim$9.58s.
 We compare the COCOLA Score (based on ”COCOLA HPS All”) with the FAD [14, 48] metric (interpreted as a sub-FAD [13, 9]), computed using CLAP [7], EnCodec [49], and VGGish.
@@
 On Ground Truth, we compute COCOLA upper bounds with real positive pairs.
-Diff-A-Riff performs best on MoisesDB, and SA ControlNet on MUSDB18-HQ.
+Diff-A-Riff performs best on MoisesDB and SA ControlNet on MUSDB18-HQ.
 V-C MOS vs COCOLA Score
 To evaluate COCOLA’s effectiveness, we conducted a subjective test with 41 participants who rated (0-4) the coherence of (pitch) Shifted and (time) Stretched chunks.
-We also used Random, Ground Truth, Diff-A-Riff and SA ControlNet tracks from the previous experiment.
+We also used Random, Ground Truth, Diff-A-Riff, and SA ControlNet tracks from the previous experiment.
 We average the ratings (MOS) for each chunk.
 Figure 3 shows the correlation plot between MOS and COCOLA Score (using ”COCOLA HPS All”).
-The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$ indicating a good correlation with high statistical signficance.
+The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$, having high statistical significance.
 VI Conclusion
-In this paper, we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems.
-Both classification experiments and human scores demonstrate the efficacy of the method.
-We used COCOLA for benchmarking music accompaniment models, proposing a new evaluation metric for the task.
-We plan to improve the quality of COCOLA by training on additional stem-level datasets [46] or on pre-separated [13] large scale datasets [43].
+In this paper, we proposed COCOLA, a contrastive encoder to evaluate the coherence between musical stems.
+Classification experiments and human scores demonstrate the effectiveness of the method.
+We used COCOLA to benchmark music accompaniment models, proposing a new evaluation metric for the task.
+We aim to improve COCOLA using additional stem-level [50] or pre-separated large-scale [47] datasets.